OpenStudy (anonymous):

What is the midpoint of (–2, 0) and (4, –2)? (0, 0) (2, –2) (1, –1) (–6, 2)

OpenStudy (anonymous):

im guessing 2,-2

hartnn (hartnn):

the mid point of points (a,b)and (c,d) is given by (m,n) where m= (a+c)/2 n= (b+d)/2 that is , u just have to take the average of respective co-ordinates

hartnn (hartnn):

nopes.

OpenStudy (anonymous):

alright

hartnn (hartnn):

so what is (4-2)/2 = ?

OpenStudy (anonymous):

Could you explain differently?

OpenStudy (anonymous):

1,-1

hartnn (hartnn):